OpenStudy (hatahoooochee):

Help, I'll fan and medal Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = Four divided by x squared. + 9

OpenStudy (hatahoooochee):

Answers: f(x) = x + 9, g(x) = Four divided by x squared. f(x) = x, g(x) = Four divided by x. + 9 f(x) = One divided by x., g(x) = Four divided by x. + 9 f(x) = Four divided by x squared., g(x) = 9

OpenStudy (hatahoooochee):

@Vocaloid @agent0smith

OpenStudy (phi):

f( g(x) ) means "replace x" with the symbols (or string) g(x) for example , if you had f(x) = 2 x +3 then f (g(x)) = 2 g(x) + 3 (notice how we erased the x, and put in g(x) instead)

OpenStudy (phi):

now if g(x) has a "definition", for example we are told g(x) = x^3 then we can go back and put in the "full definition" in place of the short name g(x)

OpenStudy (phi):

For example, say f(x) = 2 x + 3 and we do f(g(x)) = 2 g(x) + 3 now say g(x) = x^3 +3 then we can replace g(x) in f(g(x)) = 2 g(x) + 3 to get f(g(x)) = 2 ( x^3+3) + 3

OpenStudy (phi):

I would try to do that for each of your choices and see if any match up with the pattern you want.